OpenStudy (anonymous):

what is answer for: 8(4p+1)=32p+8

OpenStudy (anonymous):

32p + 8 = 32p + 8 every answer is correct

OpenStudy (turingtest):

do you see what henkjan means?

OpenStudy (anonymous):

It's not completed

OpenStudy (turingtest):

henkjan is correct, but you can illustrate the point more clearly. do you agree up to 8(4p+1)=32p+8 32p+8=32p+8 ???

OpenStudy (king):

what he means is that whatever value u put fr p it is rite.......

OpenStudy (anonymous):

thanks

OpenStudy (turingtest):

well, hankjan should hve said "all real number" or something, not "every answer is correct", that's a little silly

OpenStudy (king):

actually everuthing is correct it doesnt matter what type of number it is.......

OpenStudy (turingtest):

@king, that seems true, but why do they like to write the answer as\[\left\{ x:x \in \mathbb{R} \right\}\]can you tell me 'cause I don;t kow

OpenStudy (king):

no the number doesnt have to be only a real number it can be anything a complex number also

OpenStudy (turingtest):

I see what you're saying, but why do they say ALL REALS in class? I'm just wondering...